# Synergy: Towards On-Body AI via Tiny AI Accelerator Collaboration on Wearables

### v2

@@ 4.2. Device-Agnostic Programming Interface
4.3. Holistic Collaboration Planning
@@
An execution plan is defined as a sequence of tuples, each containing a task and an assigned device (Figure 6 (b)). Synergy supports seven task types: (i) sensing, (ii) data loading to accelerator memory, (iii) (partial) model inference, (vi) data unloading from accelerator memory, (v) Tx, (vi) Rx, and (vii) interaction. For two wearables, smart glasses and a smart ring, a pipeline can be described as (camera on glasses, EfficientNet, haptic feedback); EfficientNet has 29 layers. Then, one execution plan example would be [camera $\rightarrow$ data loading $\rightarrow$ $\textit{EfficientNet}^{0:19}$rightarrow data unloading $\rightarrow$ Tx to ring] for the glasses and [Rx from glasses $\rightarrow$ data loading $\rightarrow$ $\textit{EfficientNet}^{19:29}$rightarrow data unloading $\rightarrow$ haptic] for the smart ring, where $\textit{EfficientNet}^{i:j}$ refers to the model’s layer $i$ to $j$.
@@
Runnable holistic collaboration plan generation: The next step is to generate holistic collaboration plans, each integrating execution plans from all pipelines (one execution plan per each app). It gives Synergy visibility over resource competition and dependencies across devices (see Figure 6 (b)). A key consideration is filtering out plans that cause OOR cases. We consider three memory constraints: (i) weight memory, (ii) bias memory, and (iii) the maximum number of layers supported (e.g., 442 KB, 2 KB, 32 for MAX78000 (Max78000, [n. d.]) and 1.3 MB, 2 MB, 128 for MAX78002 (Max78002, [n. d.])). A collaboration plan is runnable if the total weight memory, bias memory, and number of layers for assigned tasks do not exceed the AI accelerators’ capacities.
4.4. Progressive Search Space Reduction
@@
We propose a progressive search space reduction method to reduce the intractable search space for the holistic collaboration plan selection. For each pipeline $p$, the number of execution plans available $N_{p}$ can be modeled as: $N_{p}=\sum_{d=1}^{D}{}_{D}\mathrm{P}_{d}\cdot{}_{L-1}\mathrm{C}_{d-1}\cdot D^{2}$, where $D$ and $L$ denote the number of available devices and model layers, respectively. ${}_{D}\mathrm{P}_{d}$ represents possible device orders to use, ${}_{L-1}\mathrm{C}_{d-1}$ is model splitting candidates, and $D^{2}$ is all source/target device mappings, respectively. This search space grows exponentially for holistic collaboration planning with concurrent pipelines, i.e., $O(N_{p_{1}}\times N_{p_{2}}\times...)$. For instance, for three small models (a 9-layer KWS, a 14-layer SimpleNet, and a 19-layer UNet in §6.1) with three MAX78000 devices, the number of holistic collaboration plans that can be generated is $1,971\times 4,941\times 9,261=90,190,202,571$. It is evident that a complete search algorithm considering all combinations requires substantial computational overhead. To mitigate this, we propose data intensity-aware execution plan accumulation, shown in Figure 7. This approach arranges execution plans in sequence via our data intensity metric and selects an execution plan for one app at a time, building on the collaboration plan integrated with previously selected execution plans from earlier apps, thereby reducing complexity to $O(N_{p_{1}}+N_{p_{2}}+...)$.
@@
Figure 7. Comparison between complete search (a) and our progressive pipeline selection (b).
@@
Pipeline prioritization: The key to data intensity-aware execution plan accumulation is determining the sequence of pipelines. Synergy prioritizes data-intensive pipelines for execution plan selection, meaning pipelines that require the transmission of larger sensing data and intermediate outputs. The intuition is that optimizing data-intensive pipelines yields greater overall benefits for the system. To understand the relationship between data intensity (i.e., data size) and latency, we conducted a preliminary measurement. Figure 8 shows layer-wise latency for inference, memory, and communication, alongside output data size using UNet (Ronneberger et al., 2015) as an example. Compared to inference latency (1.5 ms), both memory latency (10.6 ms) and communication latency (6869.1 ms) are significantly higher, being 7×7$\times$7 × and 4579×4579$\times$4579 × greater, respectively. This latency gap between inference and communication is notable in tiny AI accelerators designed for fast inference. Moreover, the output size varies significantly across different layers, with the lowest layer latency at 4426.2 ms and the highest at 161864.5 ms, indicating a 36×36$\times$36 × difference. These findings suggest that (i) the end-to-end latency of a pipeline varies greatly depending on the model split in the execution plan, (ii) latency is strongly related to the size of data exchanged, and (iii) prioritizing data-intensive pipelines could offer more optimization opportunities due to fewer resource conflicts with other pipelines.
Figure 8. Layer-wise latency analysis for UNet. The y-axis is in the logarithmic scale.
@@
Pipeline arrangement and accumulation: We define a pipeline’s data intensity as the average data size of transmission over all execution plans. Specifically, given the input size ${\tt In}^{\tt size}$ and output size ${\tt Out}^{\tt size}_{l}$ for each layer $l\in L$, the data intensity of a pipeline is defined as: $({\tt In}^{\tt size}+\sum_{l}{\tt Out}^{\tt size}_{l})/(L+1)$. The pipelines are sorted in descending order with this metric. Then, for each pipeline, Synergy generates runnable holistic collaboration plans by combining each execution plan of the current pipeline with the previously selected execution plans from earlier pipelines. Then, Synergy selects the execution plan of which holistic collaboration plan is expected to have the highest system-wide throughput based on our online throughput estimation (§4.5).
@@
Figure 9. Comparison among different pipeline prioritization strategies against complete search (Oracle).
@@
Comparison with other prioritization methods: We investigate (i) how significantly our selection reduces the search space compared to a complete search, and (ii) how close our prioritization method is to the optimal selection compared to other alternatives. We compare Synergy with seven baselines: the complete search (Oracle), ascending order of data intensity (DataIntensityAsc), descending/ascending order of model size (ModelSizeDes/ModelSizeAsc), and descending/ascending order of the number of model layers (NumLayersDes/NumLayersAsc), and without prioritization (Sequential). Except for Oracle, the six baselines use the execution plan accumulation method but with different prioritization objectives. In this experiment, we assume three concurrent pipelines with two MAX78000 devices and consider all possible combinations of three pipelines out of eight (Table 1). Figure 9 shows the relative throughput ratio compared to Oracle by averaging the throughput over all pipeline combinations. The results validate our design choice. Synergy outperforms the other alternatives, showing only a 3.9% degradation in overall throughput compared to the complete search. Additionally, the execution plan accumulation approach (Synergy and other prioritizations) results in a 5576$\times$ reduction in search space compared to Oracle.
4.5. Throughput Estimation for Distributed Tiny AI Accelerators
@@
Figure 10. Tasks involved in a pipeline for latency modeling ((1)-(6)) from the viewpoint of a device.
@@
For holistic collaboration plan selection, evaluating the throughput of a plan is necessary. To achieve this, we first model the execution latency for each task and estimate throughput based on this latency model. Figure 10 illustrates the operation of the five types of tasks ((1)-(6)) involved in a device. We present (i) our novel latency estimation model for model inference tasks, which reflects the unique characteristics of tiny AI accelerators (§4.5.1), and (ii) the latency estimation for other tasks (§4.5.2). We then introduce our technique for estimating the system-wide throughput of a holistic collaboration plan (§4.5.3).
@@
Figure 11. Correlations of latency with (a) trainable parameters (left) and (b) accelerator clock cycles (right). Each dot means a different layer within a model.
4.5.1. Latency Modeling for Tiny AI accelerators
@@
For latency estimation of model splitting options, existing studies rely on measurement-driven approaches: online latency measurement by running each model layer on the target device directly (Laskaridis et al., 2020; Zhang et al., 2020) or offline latency modeling with model configurations (Kang et al., 2017; Jeong et al., 2018; Wang et al., 2021), such as learning a regression model based on the number of model parameters and latency. The first approach—direct runtime measurement for each layer—is accurate and feasible for two-tier architectures (e.g., a smartphone and a cloud server) with limited cases but entails significant overhead in our scenario involving multiple models and resource-constrained devices. The second approach—latency modeling with offline profiling—mitigates this issue but has critical limitations in our environment. First, its scalability is highly limited, requiring profiling of all possible models and AI accelerators before app deployment. Second, due to the unique characteristics of AI accelerators, existing configuration-based latency modeling may not work. Figure 11 (a) shows the relationship between the number of model parameters and the corresponding inference latency of each layer in eight models on MAX78000. Due to hardware-level optimization in tiny AI accelerators, there is a weak correlation between these two variables.
@@
(1) $\mathcal{L}_{\tt ai\_acc}=\mathcal{L}_{\tt load}+\mathcal{L}_{\tt inf}+\mathcal{L}_{\tt unload},$
@@
(2) ^$\hat{\mathbb{C}}_{\tt MLP}$ $=H_{\tt in}\cdot W_{\tt in}\cdot C_{\tt in}\cdot C_{\tt out}$
(3) ^$\hat{\mathbb{C}}_{\tt CNN}$ $=K^{2}\cdot H_{\tt in}\cdot W_{\tt out}\cdot C_{\tt in}\cdot C_{\tt out}$
@@
(4) $\mathbb{C}_{\tt MLP}$ $=H_{\tt in}\cdot W_{\tt in}\cdot\left\lceil\frac{C_{\tt in}}{P}\right\rceil\cdot C_{\tt out}$
(5) $\mathbb{C}_{\tt CNN}$ $=H_{\tt in}\cdot W_{\tt out}\cdot\left\lceil\frac{C_{\tt in}}{P}\right\rceil\cdot C_{\tt out}$
@@
4.5.2. Latency Modeling for Other Tasks
Besides operations on tiny AI accelerators, on-body AI apps involve sensing and communication (Rx and Tx) tasks ((1), (5), (6) in Figure 11). For sensing tasks, we measure the latency for camera and audio inputs with different parameters (e.g., different sampling rates) during the profiling phase and estimate the latency by matching the app’s requirements to the profiles. For communication latency, similar to memory operations on tiny AI accelerators, we estimate it by dividing the data size by the bandwidth of the wireless channel. Advanced methods (Kang et al., 2017; Zhang et al., 2020) can address wireless transmission latency fluctuations, but this is beyond the scope of this paper.
4.5.3. Throughput Estimation of Holistic Collaboration Plan
@@
Figure 12. Comparison between the conventional approach (a) and our proposed adaptive parallelization mechanisms (b, c) across two devices ($d_{1}$ and $d_{2}$).
4.6. Adaptive Task Parallelization
After selecting the best holistic collaboration plan, Synergy deploys the tasks and their connection to the corresponding devices. As existing model partitioning techniques (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021) focused on a single run for a single model, a straightforward way to execute a holistic collaboration plan with concurrent pipelines over multiple runs (continuous inference over time) is to run each pipeline sequentially—immediately run another after the completion of the preceding pipeline—in a continuous manner, as shown in Figure 12 (a). However, since tasks in the plan runs over different computation units (processors, AI accelerator, and communication module) from distributed devices, such a sequential execution yields long idle time for unused units. The challenge is how to enhance performance by effectively utilizing these diverse computation units while keeping the sequential dependency of tasks within each pipeline.
@@ 4.6. Adaptive Task Parallelization
5. Prototype Implementation
We prototyped Synergy on the off-the-shelf MAX78000 feather board (Max78000FTHR, [n. d.]) and MAX78002 Evaluation Kit (Max78002EVKIT, [n. d.]), which are a development platform for the MAX78000 (Max78000, [n. d.]) and MAX78002 (Max78002, [n. d.]), respectively (Figure 13). Note that although the development platform is bulky, the actual size of the accelerators is tiny (e.g., 8mm×$\times$8mm for MAX78000).
Figure 13. Hardware setup.
For wireless communication, we interfaced MAX78000/ MAX78002 with an ESP8266 Wi-Fi module (ESP8266, [n. d.]). To ensure that the communication process via Wi-Fi does not disrupt the data transmissions with MAX78000/MAX78002, we have employed a round-robin scheduling algorithm on ESP8266. This mechanism is designed to alternate the data flow: transmitting data to the wireless communication channel for outward transmission and redirecting to the serial communication channel upon receiving data from the Wi-Fi interface. Additionally, we employ lightweight Internet Protocol a compact, standalone implementation of the TCP/IP protocol tailored for low-power devices, consuming a mere 40KB of memory.

### v3

@@ IV-B Device-Agnostic Programming Interface
IV-C Holistic Collaboration Planning
@@
An execution plan is defined as a sequence of tuples, each containing a task and an assigned device (Figure 6 (b)). Synergy supports seven task types: (i) sensing, (ii) data loading to accelerator memory, (iii) (partial) model inference, (vi) data unloading from accelerator memory, (v) Tx (transmitting data), (vi) Rx (receiving data), and (vii) interaction. For two wearables, smart glasses and a smart ring, a pipeline can be described as (camera on glasses, EfficientNet, haptic feedback); EfficientNet has 29 layers. Then, one execution plan example would be [camera $\rightarrow$ data loading $\rightarrow$ $\textit{EfficientNet}^{0:19}$rightarrow data unloading $\rightarrow$ Tx to ring] for the glasses and [Rx from glasses $\rightarrow$ data loading $\rightarrow$ $\textit{EfficientNet}^{19:29}$rightarrow data unloading $\rightarrow$ haptic] for the smart ring, where $\textit{EfficientNet}^{i:j}$ refers to the model’s layer $i$ to $j$.
@@
Runnable holistic collaboration plan generation: The next step is to generate holistic collaboration plans, each integrating execution plans from all pipelines (one execution plan per each app). It gives Synergy visibility over resource competition and dependencies across devices (see Figure 6 (b)). A key consideration is filtering out plans that cause OOR cases. We consider three memory constraints: (i) weight memory, (ii) bias memory, and (iii) the maximum number of layers supported (e.g., 442 KB, 2 KB, 32 for MAX78000 [1] and 1.3 MB, 2 MB, 128 for MAX78002 [2]). A collaboration plan is runnable if the total weight memory, bias memory, and number of layers for assigned tasks do not exceed the AI accelerators’ capacities.
IV-D Progressive Search Space Reduction
@@
We propose a progressive search space reduction method to reduce the intractable search space for the holistic collaboration plan selection. For each pipeline $p$, the number of execution plans available $N_{p}$ can be modeled as: $N_{p}=\sum_{d=1}^{D}{}_{D}\mathrm{P}_{d}\cdot{}_{L-1}\mathrm{C}_{d-1}\cdot D^{2}$, where $D$ and $L$ denote the number of available devices and model layers, respectively. ${}_{D}\mathrm{P}_{d}$ represents $d$-permutation of $D$, which means possible device orders to use, ${}_{L-1}\mathrm{C}_{d-1}$ is $d-1$ combinations of $L-1$, which is model splitting candidates, and $D^{2}$ is all source/target device mappings, respectively. This search space grows exponentially for holistic collaboration planning with concurrent pipelines, i.e., $O(N_{p_{1}}\times N_{p_{2}}\times...)$. For instance, for three small models (a 9-layer KWS, a 14-layer SimpleNet, and a 19-layer UNet in §VI-A) with three MAX78000 devices, the number of holistic collaboration plans that can be generated is $1,971\times 4,941\times 9,261=90,190,202,571$. It is evident that a complete search algorithm considering all combinations requires substantial computational overhead. To mitigate this, we propose data intensity-aware execution plan accumulation, shown in Figure 7. This approach arranges execution plans in sequence via our data intensity metric and selects an execution plan for one app at a time, building on the collaboration plan integrated with previously selected execution plans from earlier apps, thereby reducing complexity to $O(N_{p_{1}}+N_{p_{2}}+...)$.
@@
Figure 7: Comparison between complete search (a) and our progressive pipeline selection (b).
@@
Pipeline prioritization: The key to data intensity-aware execution plan accumulation is determining the sequence of pipelines. Synergy prioritizes data-intensive pipelines for execution plan selection, meaning pipelines that require the transmission of larger sensing data and intermediate outputs. The intuition is that optimizing data-intensive pipelines yields greater overall benefits for the system. To understand the relationship between data intensity (i.e., data size) and latency, we conducted a preliminary measurement. Figure 8 shows layer-wise latency for inference, memory, and communication, alongside output data size using UNet [31] as an example. Compared to inference latency (1.5 ms), both memory latency (10.6 ms) and communication latency (6869.1 ms) are significantly higher, being 7×7$\times$7 × and 4579×4579$\times$4579 × greater, respectively. This latency gap between inference and communication is notable in tiny AI accelerators designed for fast inference. Moreover, the output size varies significantly across different layers, with the lowest layer latency at 4426.2 ms and the highest at 161864.5 ms, indicating a 36×36$\times$36 × difference. These findings suggest that (i) the end-to-end latency of a pipeline varies greatly depending on the model split in the execution plan, (ii) latency is strongly related to the size of data exchanged, and (iii) prioritizing data-intensive pipelines could offer more optimization opportunities due to fewer resource conflicts with other pipelines.
Figure 8: Layer-wise latency analysis for UNet. The y-axis is in the logarithmic scale. Since the scale is logarithmic, the total height of each stacked bar represents the geometric sum of its components, making small values appear relatively smaller while highlighting differences across orders of magnitude.
@@
Pipeline arrangement and accumulation: We define a pipeline’s data intensity as the average data size of transmission over all execution plans. Specifically, given the input size ${\tt In}^{\tt size}$ and output size ${\tt Out}^{\tt size}_{l}$ for each layer $l\in L$, the data intensity of a pipeline is defined as: $({\tt In}^{\tt size}+\sum_{l}{\tt Out}^{\tt size}_{l})/(L+1)$. The pipelines are sorted in descending order with this metric. Then, for each pipeline, Synergy generates runnable holistic collaboration plans by combining each execution plan of the current pipeline with the previously selected execution plans from earlier pipelines. Then, Synergy selects the execution plan of which holistic collaboration plan is expected to have the highest system-wide throughput based on our online throughput estimation (§IV-E).
@@
Figure 9: Comparison among different pipeline prioritization strategies against complete search (Oracle).
@@
Comparison with other prioritization methods: We investigate (i) how significantly our selection reduces the search space compared to a complete search, and (ii) how close our prioritization method is to the optimal selection compared to other alternatives. We compare Synergy with seven baselines: the complete search (Oracle), ascending order of data intensity (DataIntensityAsc), descending/ascending order of model size (ModelSizeDes/ModelSizeAsc), and descending/ascending order of the number of model layers (NumLayersDes/NumLayersAsc), and without prioritization (Sequential). Except for Oracle, the six baselines use the execution plan accumulation method but with different prioritization objectives. In this experiment, we assume three concurrent pipelines with two MAX78000 devices and consider all possible combinations of three pipelines out of eight (Table I). Figure 9 shows the relative throughput ratio compared to Oracle by averaging the throughput over all pipeline combinations. The results validate our design choice. Synergy outperforms the other alternatives, showing only a 3.9% degradation in overall throughput compared to the complete search. Additionally, our execution plan accumulation approach results in a 5576$\times$ reduction in search space compared to Oracle. Note that the other prioritization methods have the same search space as ours because they are all based on our progressive pipeline selection.
IV-E Throughput Estimation for Distributed Tiny AI Accelerators
@@
Figure 10: Tasks involved in a pipeline for latency modeling ((1)-(6)) from the viewpoint of a device.
@@
For holistic collaboration plan selection, evaluating the throughput of a plan is necessary. To achieve this, we first model the execution latency for each task and estimate throughput based on this latency model. Figure 10 illustrates the operation of the five types of tasks ((1)-(6)) involved in a device. We present (i) our novel latency estimation model for model inference tasks, which reflects the unique characteristics of tiny AI accelerators (§IV-E1), and (ii) the latency estimation for other tasks (§IV-E2). We then introduce our technique for estimating the system-wide throughput of a holistic collaboration plan (§IV-E3).
@@
Figure 11: Correlations of latency with (a) trainable parameters (left) and (b) accelerator clock cycles (right). Each dot means a different layer within a model.
IV-E1 Latency Modeling for Tiny AI accelerators
@@
For latency estimation of model splitting options, existing studies rely on measurement-driven approaches: online latency measurement by running each model layer on the target device directly [13, 12] or offline latency modeling with model configurations [10, 14, 15], such as learning a regression model based on the number of model parameters and latency. The first approach—direct runtime measurement for each layer—is accurate and feasible for two-tier architectures (e.g., a smartphone and a cloud server) with limited cases but entails significant overhead in our scenario involving multiple models and resource-constrained devices. The second approach—latency modeling with offline profiling—mitigates this issue but has critical limitations in our environment. First, its scalability is highly limited, requiring profiling of all possible models and AI accelerators before app deployment. Second, due to the unique characteristics of AI accelerators, existing configuration-based latency modeling may not work. Figure 11 (a) shows the relationship between the number of model parameters and the corresponding inference latency of each layer in eight models on MAX78000. Due to hardware-level optimization in tiny AI accelerators, there is a weak correlation between these two variables.
@@
$\mathcal{L}_{\tt ai\_acc}=\mathcal{L}_{\tt load}+\mathcal{L}_{\tt inf}+\mathcal{L}_{\tt unload},$ (1)
@@
^$\hat{\mathbb{C}}_{\tt MLP}$ $=H_{\tt in}\cdot W_{\tt in}\cdot C_{\tt in}\cdot C_{\tt out}$ (2)
^$\hat{\mathbb{C}}_{\tt CNN}$ $=K^{2}\cdot H_{\tt in}\cdot W_{\tt out}\cdot C_{\tt in}\cdot C_{\tt out}$ (3)
@@
$\mathbb{C}_{\tt MLP}$ $=H_{\tt in}\cdot W_{\tt in}\cdot\left\lceil\frac{C_{\tt in}}{P}\right\rceil\cdot C_{\tt out}$ (4)
$\mathbb{C}_{\tt CNN}$ $=H_{\tt in}\cdot W_{\tt out}\cdot\left\lceil\frac{C_{\tt in}}{P}\right\rceil\cdot C_{\tt out}$ (5)
@@
IV-E2 Latency Modeling for Other Tasks
Besides operations on tiny AI accelerators, on-body AI apps involve sensing and communication (Rx and Tx) tasks ((1), (5), (6) in Figure 11). For sensing tasks, we measure the latency for camera and audio inputs with different parameters (e.g., different sampling rates) during the profiling phase and estimate the latency by matching the app’s requirements to the profiles. For communication latency, similar to memory operations on tiny AI accelerators, we estimate it by dividing the data size by the bandwidth of the wireless channel. Advanced methods [10, 12] can address wireless transmission latency fluctuations, but this is beyond the scope of this paper.
IV-E3 Throughput Estimation of Holistic Collaboration Plan
@@
Figure 12: Comparison between the conventional approach (a) and our proposed adaptive parallelization mechanisms (b, c) across two devices ($d_{1}$ and $d_{2}$).
IV-F Adaptive Task Parallelization
After selecting the best holistic collaboration plan, Synergy deploys the tasks and their connection to the corresponding devices. As existing model partitioning techniques [10, 11, 12, 13, 14, 15, 16] focused on a single run for a single model, a straightforward way to execute a holistic collaboration plan with concurrent pipelines over multiple runs (continuous inference over time) is to run each pipeline sequentially—immediately run another after the completion of the preceding pipeline—in a continuous manner, as shown in Figure 12 (a). However, since tasks in the plan runs over different computation units (processors, AI accelerator, and communication module) from distributed devices, such a sequential execution yields long idle time for unused units. The challenge is how to enhance performance by effectively utilizing these diverse computation units while keeping the sequential dependency of tasks within each pipeline.
@@ IV-F Adaptive Task Parallelization
V Prototype Implementation
We prototyped Synergy on the off-the-shelf MAX78000 feather board [32] and MAX78002 Evaluation Kit [33], which are a development platform for the MAX78000 [1] and MAX78002 [2], respectively (Figure 13). Note that although the development platform is bulky, the actual size of the accelerators is tiny (e.g., 8mm×$\times$8mm for MAX78000).
Figure 13: Hardware setup.
For wireless communication, we interfaced MAX78000/ MAX78002 with an ESP8266 Wi-Fi module [34]. To ensure that the communication process via Wi-Fi does not disrupt the data transmissions with MAX78000/MAX78002, we have employed a round-robin scheduling algorithm on ESP8266. This mechanism is designed to alternate the data flow: transmitting data to the wireless communication channel for outward transmission and redirecting to the serial communication channel upon receiving data from the Wi-Fi interface. Additionally, we employ lightweight Internet Protocol a compact, standalone implementation of the TCP/IP protocol tailored for low-power devices, consuming a mere 40KB of memory.